Verizon LTE 10:13 AM 66% Math 2193.docx Done Preface: I am an adjunct dental hygiene instructor in the Dental Hygiene Program at Southwestern College. The current senior class has 30 dental hygiene students graduating at the end of this month. The 30 senior students have treated patients during their second semester (8 patients required), third semester (12 patients required) and fourth semester (18 patients required) during the program. The total number of patients treated in the dental hygiene clinic were 1,140 1. The population of the study were patients treated at the no- cost dental hygiene clinic at Southwestern College. 2. The Dental Hygiene Program was started at Southwestern College in 1999. The two primary goals of establishing the program were to make dental hygiene education available in San Diego county and to provide free dental hygiene services to the minority populations that surround the clinic. The second goal is the focus of this sample. One of the Dental Hygiene Program survey questions addresses the population from which the patient identifies with (Hispanic, Black, Caucasian, Asian, American Indian, Pacific Islander, Other). 3. The sampling method used is an in-person, written survey completed by each patient at the end of his/her treatment. 4. The hypothesis to be tested is: Due to the location of the Southwestern College Dental Hygiene Program Clinic, the minority populations in the community are receiving free dental hygiene treatment. S. Type of test: one population-the patients receiving treatment in the Southwestern College Dental Hygiene Program Clinic. 6. The parameter to be tested The statistic to be tested 7. The sample size ,140 patients
Answer these questions: 1. Identify a population to study. 2.Identify the sampling method to be utilized and why 3. How is the sample to be collected? (survey, in person questions, email, facebook, etc). 4. Identify the Hypotheses to be tested. s. Identify the type of test (one population, two population). 6.Identify the Parameter and Statistic to be tested (mean, proportion, etc). 7. Collect Data on n230, state sample size: 8. Follow the steps to Hypotheses testing, analyze your data using statistical software. Include any applicable graphs and descriptive statistics 9Summarize your findings in a paragraph 4) Describe the relationship between the variables. (eg, The jumping distance is predicted to increase by... for each additional inch of height...) 5) Verify that your model is appropriate (or take remedial action if needed) by checking the assumptions of the model (e.g Based on this plot the error variance appeared non-constant, so . Specifics for the Project 1) If the data come from a sample: Define the desired target population and describe how the sample was collected. If you were not able to sample from the desired population, state what differences you might expect between the population that was actually sampled from and the desired target population. If you were not able to take a simple random sample from the population, discuss how the sampling could be improved if you were allowed more money and time If the data come from an experiment: Describe how the experiment was carried out, describe any sources of extra variation (e.g. changing temperature, different people making the measurements, etc...). Did you try to control these? Discuss how the experiment could be improved if you were allowed (more) money and time. 2) Statistics: The report of the test you performed should include at least the following statistics: a confidence interval, the p- value for testing, and descriptive statistics & graphs. Make sure and tell the reader why these statistics should be useful to them, and interpret them in the context of your data set
Answer #1

Test statistic Z since the sample size is > 30(large sample)

sample size n = 1140

X = number of success = 38

prob of success p= 8/1140+12/1140 +18/1140 =0.034

null hypothesis H0 = Due to location of the college the minority population in the community are not receiving free dental hyg treatment

Z = A-np = 38-1140 × .034 νΛ 1140 × .034 × .97 = 0.12

since |z| < 3 null hypothesis accepted